OpenStudy (anonymous):

a ball is fired vertically into the air. Its height in feet t seconds later in feet t seconds later is given by the formula: h(t)= -2t^2+80t a) what is the height of the ball after 3 sec? b) what is the velocity of the ball after 5 sec? c) when does the ball reach its man height? d) what is the max height the ball reaches?

OpenStudy (anonymous):

For a and b, just substitute the value of t. a. h(3) = -2 (3^2) + 80 (3) h(3)= -2(9) + 240 h(3) = -18 + 240 h(3) = 222 So, the ball is 222 feet from the ground. You should be able to solve for letter b now :) For c, you should use the formula -b/2a. For d, you should use the formula c^2(h)(a)

OpenStudy (anonymous):

Given the equation h(t) = -2t^2 +80t + 0 wherein -2= a, 80= b and 0 =c

OpenStudy (anonymous):

** For d, its c^2-h(a) My bad :D